OpenStudy (anonymous):

What is the solution to the system of equations? y=2x+6 3x-y=6

OpenStudy (alexandervonhumboldt2):

substitute 2x+6 for y into second equation and solve for x

OpenStudy (alexandervonhumboldt2):

y=2x+6 3x-y=6 substitute: 3x-(2x+6)=6 solve for x

OpenStudy (anonymous):

12?

OpenStudy (phi):

yes

OpenStudy (anonymous):

so 12 is the answer?

OpenStudy (phi):

x is 12. you also have to find y use y=2x+6 and replace x with 12

OpenStudy (anonymous):

y=30?

OpenStudy (phi):

Alex used "substitution" for this problem. If we use "eliminattion" we would write both equations in standard form: y=2x+6 becomes -2x + y = 6 now add the two equations -2x + y = 6 3x - y= 6 x +0 = 12 or x=12 and for y: -2x+y= 6 with x=12 we get -2*12 + y= 6 -24+y= 6 add 24 to both sides y= 30

OpenStudy (phi):

the answer is (12,30)

OpenStudy (anonymous):

Thank you so much! :)